4. Suppose that your production function is q(L) = 10L^1/2

a. In a table, calculate the output, average product and marginal product for L = 1,2,3,4,5,6,7,8,9.

b. Suppose fixed costs are $100 and the price of labor is $20. In a table, calculate the average cost and marginal cost for L = 1,2,3,4,5,6,7,8,9.

c. If you could sell your output for $9 how many workers would you hire?

Explain all answers
Answer #1

q = 10L1/2

(a) AP = q/L = (10L1/2)/L = 10L-1/2

MP = dq/dL = 1/2(10)L-1/2 = 5L-1/2

(b) Fixed cost = $100

w = $20

TC = FC+ wL

AC= TC/q

MC = (Change in TC/Change in q)

L q AP=q/L MP= (Change in TP/Change in L) VMPL=(P)(MPL)=($9)(MPL) TC= FC+wL AC=TC/q MC= (Change in TC/Change in q)
1 10 10 5 45 120 12 -
2 14.14 7.07 3.54 31.86 140 9.90 4.83
3 17.32 5.77 2.89 26.01 160 9.24 6.29
4 20 5 2.5 22.5 180 9 7.46
5 22.36 4.47 2.24 20.16 200 8.94 8.47
6 24.49 4.08 2.04 18.36 220 8.98 9.38
7 26.46 3.78 1.89 17.01 240 9.07 10.15
8 28.28 3.53 1.77 15.93 260 9.19 10.98
9 30 3.33 1.67 15.03 280 9.33 11.63

(c) P= $9

Optimal labor input is where VMPL = wage.

Value of marginal product of labor , VMPL = (P)(MPL)

We can see that when L=5 , VMPL >Wage

and when L=6 , VMPL<Wage.

Therefore, if price =$9 , then the firm would hire 5 workers.